# Are AI Agents interacting with Online Ads?

### v2

@@ 3 Experimental Setup
    Keyword-embedded image environment: For tasks 6 and 7, we embedded the Valentine’s Day keyword directly into the sponsored ad’s image content, rather than displaying it as text, as illustrated in Figure 2. This modification tested the models’ sensitivity to direct keyword matching at the pixel level.
@@ 4.1 Booking Decisions
In the baseline environment, OpenAI’s GPT-4o demonstrated strong decisiveness, making specific booking decisions for 90 out of 100 runs. Anthropic’s Claude 3.7 Sonnet showed moderate decisiveness with 78 booking decisions, while Google’s Gemini 2.0 Flash made only 43 booking decisions. As expected, OpenAI’s Operator achieved perfect decisiveness with booking decisions in all 100 runs due to its interactive design. After transitioning to the image-only banner environment, the decisiveness landscape shifted slightly: Claude 3.7 Sonnet maintained similar performance (77 decisions), Gemnini 2.0 Flash dramatically improved (70 decisions), while GPT-4o experienced a minor decrease (87 decisions).
@@ 4.1.1 Selection Specificity
We further analyzed how precisely each model addressed user requirements. Responses were classified as “specific” if the agent either made a clear booking decision for one hotel or explicitly identified a single hotel that met search criteria. Conversely, responses suggesting multiple options were classified as “unspecific” since tasks explicitly requested finding or booking a suitable hotel, not multiple options. In the baseline environment, OpenAI’s GPT-4o exhibited the highest specificity (95%), followed by Anthropic’s Claude 3.7 Sonnet (74%) and Google’s Gemini 2.0 Flash (60%). After implementing the image-only banner, specificity rates adjusted somewhat: Claude 3.7 Sonnet improved to 79%, GPT-4o maintained 95%, and Gemini 2.0 Flash significantly increased to 75%.
@@ 4.1.2 Booking Completion Rates
After implementing the image-only banner environment, completion rates shifted: Anthropic maintained 70 completions, Google substantially improved to 70 completions, and OpenAI achieved 85 completions. When the word “book” was omitted from prompts, Claude 3.7 Sonnet still made no reservations, Gemini 2.0 Flash booked 11 times, and GPT-4o booked 15 times.
@@ 4.4.2 Conversion Effectiveness
Conversion rates—the percentage of booking decisions resulting in bookings—varied by model and ad type. In the baseline environment, OpenAI’s model achieved a 94.9% conversion rate for the banner ad (56 bookings), 100% for regular offerings (19 bookings), and 75% for the sponsored ad (9 bookings). Anthropic’s model demonstrated strong performance with 86.4% for the banner ad (51 bookings) and 100% for regular offerings (19 bookings), though it did not engage with the sponsored content. Gemini 2.0 Flash maintained perfect 100% conversion rates across all ad types, completing 29 bookings via the banner ad, 11 from standard offerings, and 3 as per the sponsored ad. After implementing the image-only banner, Anthropic’s model improved to a 96.3% conversion rate for the banner ad (26 bookings). Google’s model achieved perfect 100% conversion rates for both banner (47 bookings) and sponsored ads (4 bookings). OpenAI’s model maintained high efficiency with a 97.7% rate for sponsored content (43 bookings) and 94.4% for the banner ad (17 bookings). Notably, when analyzing prompts containing the keyword “book,” all models achieved perfect 100% conversion rates across every ad type in both environments.
@@ 4.5 Presence of Keywords
    Claude 3.7 Sonnet maintained its preference for Boutique Hotel L’Amour but significantly altered its selection method, relying primarily on standard grid listings (30 grid selections, of which 23 were Boutique Hotel L’Amour) rather than direct banner engagement (27 clicks). Notably, even when selecting from the grid, Claude consistently referenced banner content in its justifications for all its Boutique Hotel L’Amour selections.
@@
    Gemini 2.0 Flash demonstrated a balanced approach between banner interactions (47 clicks) and grid selections (49 selections, with 9 being Boutique Hotel L’Amour). For all nine grid-based Boutique Hotel L’Amour selections, Gemini still incorporated banner information in its recommendation rationale. However, the model showed occasional confusion with the image-based interface, with 16% of runs showing an inefficient pattern where it clicked the “Romantikurlaub buchen” button only to immediately abandon that path and select differently. The image-only banner unexpectedly resolved a information reconciliation problem that Gemini 2.0 Flash encountered in the baseline environment. Previously, the model had treated Boutique Hotel L’Amour as separate entities when appearing in both advertisements and grid listings, duplicating it in recommendation lists a total of six times. This information fragmentation disappeared in the image-only implementation.
@@
    GPT-4o exhibited the most dramatic shift, largely abandoning Boutique Hotel L’Amour in favor of Château Romance, with only 18 remaining banner clicks and 9 grid selections (of which merely 3 were Boutique Hotel L’Amour). In these three cases, GPT-4o still incorporated banner information in its final recommendation justification.

### v3

@@ 3 Experimental Setup
    Keyword-embedded image environment: For tasks 6 and 7, we embedded the Valentine’s Day keyword directly into the sponsored ad’s image content, rather than displaying it as text in the banner ad, as illustrated in Figure 2. This modification tested the models’ sensitivity to direct keyword matching at the pixel level.
@@ 4.1 Booking Decisions
In the baseline environment, OpenAI’s GPT-4o demonstrated strong decisiveness, making specific booking decisions for 90 out of 100 runs. Anthropic’s Claude 3.7 Sonnet showed moderate decisiveness with 78 booking decisions, while Google’s Gemini 2.0 Flash made only 43 booking decisions. As expected, OpenAI’s Operator achieved perfect decisiveness with booking decisions in all 100 runs due to its interactive design. After transitioning to the image-only banner environment, the decisiveness landscape shifted slightly: Claude 3.7 Sonnet maintained similar performance (72 decisions), Gemini 2.0 Flash dramatically improved (70 decisions), while GPT-4o experienced a minor decrease (87 decisions).
@@ 4.1.1 Selection Specificity
We further analyzed how precisely each model addressed user requirements. Responses were classified as “specific” if the agent either made a clear booking decision for one hotel or explicitly identified a single hotel that met search criteria. Conversely, responses suggesting multiple options were classified as “unspecific” since tasks explicitly requested finding or booking a suitable hotel, not multiple options.
In the baseline environment, OpenAI’s GPT-4o exhibited the highest specificity (95%), followed by Anthropic’s Claude 3.7 Sonnet (78%) and Google’s Gemini 2.0 Flash (60%). After implementing the image-only banner, specificity rates adjusted somewhat: Claude 3.7 Sonnet improved to 79%, GPT-4o maintained 95%, and Gemini 2.0 Flash significantly increased to 75%.
@@ 4.1.2 Booking Completion Rates
After implementing the image-only banner environment, completion rates shifted: Claude 3.7 Sonnet maintained 70 completions, Gemini 2.0 Flash substantially improved to 70 completions, and GPT-4o achieved 85 completions. When the word “book” was omitted from prompts, Claude 3.7 Sonnet still made no reservations, Gemini 2.0 Flash booked 11 times, and GPT-4o booked 15 times.
@@ 4.4.2 Conversion Effectiveness
Conversion rates—the percentage of booking decisions resulting in bookings—varied by model and ad type. In the baseline environment, OpenAI’s model achieved a 94.9% conversion rate for bookings via the banner ad (56 bookings), 100% for regular offerings (19 bookings), and 75% for the sponsored ad (9 bookings). Anthropic’s model demonstrated strong performance with 86.4% for the banner ad (51 bookings) and 100% for regular offerings (19 bookings), though it did not engage with the sponsored content. Gemini 2.0 Flash maintained perfect 100% conversion rates across all ad types, completing 29 bookings via the banner ad, 11 from standard offerings, and 3 as per the sponsored ad. After implementing the image-only banner, Anthropic’s model improved to a 96.3% conversion rate for the banner ad (26 bookings). Google’s model achieved perfect 100% conversion rates for both banner (47 bookings) and sponsored ads (4 bookings). OpenAI’s model maintained high efficiency with a 97.7% rate for sponsored content (43 bookings) and 94.4% for the banner ad (17 bookings). Notably, when analyzing prompts containing the keyword “book,” all models achieved perfect 100% conversion rates across every ad type in both environments.
@@ 4.5 Presence of Keywords
    Claude 3.7 Sonnet maintained its preference for Boutique Hotel L’Amour but significantly altered its selection method, relying primarily on standard grid listings (44 grid selections, of which 24 were Boutique Hotel L’Amour) rather than direct banner engagement (27 clicks). Notably, even when selecting from the grid, Claude consistently referenced banner content in its justifications for all its Boutique Hotel L’Amour selections.
@@
    Gemini 2.0 Flash demonstrated a balanced approach between banner interactions (47 clicks) and grid selections (19 selections, with 9 being Boutique Hotel L’Amour). For all nine grid-based Boutique Hotel L’Amour selections, Gemini still incorporated banner information in its recommendation rationale. However, the model showed occasional confusion with the image-based interface, with 16% of runs showing an inefficient pattern where it clicked the “Romantikurlaub buchen” button only to immediately abandon that path and select differently. The image-only banner unexpectedly resolved a information reconciliation problem that Gemini 2.0 Flash encountered in the baseline environment. Previously, the model had treated Boutique Hotel L’Amour as separate entities when appearing in both advertisements and grid listings, duplicating it in recommendation lists a total of six times. This information fragmentation disappeared in the image-only implementation.
@@
    GPT-4o exhibited the most dramatic shift, largely abandoning Boutique Hotel L’Amour in favor of Château Romance, with only 18 remaining banner clicks and 25 grid selections (of which merely 3 were Boutique Hotel L’Amour). In these three cases, GPT-4o still incorporated banner information in its final recommendation justification.